# Investigating photometric and spectroscopic variability in the multiply-imaged ‘Little Red Dot’ A2744-QSO1

### v2

@@ Acknowledgements.
We would like to thank Xihan Ji, Hannah Übler, and Roberto Maiolino, for cordial and useful discussions. The BGU lensing group acknowledges support by grant No. 2020750 from the United States-Israel Binational Science Foundation (BSF) and grant No. 2109066 from the United States National Science Foundation (NSF), and by the Israel Science Foundation Grant No. 864/23. P.D. warmly thanks the European Commission’s and University of Groningen’s CO-FUND Rosalind Franklin program. This work is based on observations obtained with the NASA/ESA/CSA JWST, namely programs GO-2756, -2561, -2883, -3538, -4111, and -3516, retrieved from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI). STScI is operated by the Association of Universities for Research in Astronomy, Inc. under NASA contract NAS 5-26555. The spectroscopy products presented herein, from JWST program GO-2561, were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. This work also makes use of the Center for Computational Astrophysics at the Flatiron Institute which is supported by the Simons Foundation. Support for JWST programs GO-2561, -4111, and -3516 was provided by NASA through grants from STScI. This research made use of Astropy,444http://www.astropy.org a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018) and Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2024), as well as the packages NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), and the MAAT Astronomy and Astrophysics tools for MATLAB (Ofek, 2014).References
@@
- Baggen et al. (2024) Baggen, J. F. W., van Dokkum, P., Brammer, G., et al. 2024, arXiv e-prints, arXiv:2408.07745
@@
- Bezanson et al. (2022) Bezanson, R., Labbé, I., Whitaker, K. E., et al. 2022, arXiv e-prints, arXiv:2212.04026
@@
- de Graaff et al. (2024) de Graaff, A., Brammer, G., Weibel, A., et al. 2024, arXiv e-prints, arXiv:2409.05948
- de Graaff et al. (2023) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2023, arXiv e-prints, arXiv:2308.09742
@@
- Golubchik et al. (2024) Golubchik, M., Steinhardt, C. L., Zitrin, A., et al. 2024, arXiv e-prints, arXiv:2408.00073
@@
- Inayoshi & Maiolino (2024) Inayoshi, K. & Maiolino, R. 2024, arXiv e-prints, arXiv:2409.07805
@@
- Killi et al. (2023) Killi, M., Watson, D., Brammer, G., et al. 2023, arXiv e-prints, arXiv:2312.03065
@@
- Li et al. (2024) Li, Z., Inayoshi, K., Chen, K., Ichikawa, K., & Ho, L. C. 2024, arXiv e-prints, arXiv:2407.10760
- Lin et al. (2024) Lin, R., Zheng, Z.-Y., Jiang, C., et al. 2024, arXiv e-prints, arXiv:2412.08396
@@
- Ma et al. (2024) Ma, Y., Greene, J. E., Setton, D. J., et al. 2024, arXiv e-prints, arXiv:2410.06257
@@
- Maiolino et al. (2024) Maiolino, R., Risaliti, G., Signorini, M., et al. 2024, arXiv e-prints, arXiv:2405.00504
@@
- Perger et al. (2024) Perger, K., Fogasy, J., Frey, S., & Gabányi, K. É. 2024, arXiv e-prints, arXiv:2411.19518
- Price et al. (2024) Price, S. H., Bezanson, R., Labbé, I., et al. 2024, arXiv e-prints, arXiv:2408.03920
@@
- Suess et al. (2024) Suess, K. A., Weaver, J. R., Price, S. H., et al. 2024, arXiv e-prints, arXiv:2404.13132
- Tee et al. (2024) Tee, W. L., Fan, X., Wang, F., & Yang, J. 2024, arXiv e-prints, arXiv:2412.05242
- Treu et al. (2022) Treu, T., Roberts-Borsani, G., Bradac, M., et al. 2022, arXiv e-prints, arXiv:2206.07978
@@
- Yue et al. (2024) Yue, M., Eilers, A.-C., Ananna, T. T., et al. 2024, arXiv e-prints, arXiv:2404.13290
@@ Appendix B Spectroscopic measurements
For spectroscopy, we use the latest reductions of the UNCOVER spectroscopic data (Price et al. 2024) available from the DJA (Heintz et al. 2024; de Graaff et al. 2024). The spectra are reduced with MSAEXP v0.9.5 and reach m, fully covering H at . The spectra of images A and B have typical per pixel in the continuum, while the spectrum of images C, and A in epoch 8, have . In addition to the uncertainties provided by MSAEXP, we adopt a systematic error floor of 3 % per pixel in the continuum and 6 % per pixel at the location of strong lines. These errors are added in quadrature before fitting the spectra to account for spectro-photometric calibration uncertainties and other uncertainties related to the detailed shape of the intrinsic lines and the NIRSpec line-spread-function (LSF). The spectra are modeled over an observed wavelength-range m, i.e. red-ward of the red-shifted Balmer break, as we are here primarily interested in a robust measurement of the strongest Balmer lines and the continuum at the location of said lines.
@@
The model fitted to the spectra consists of the following components: The continuum is modeled with a power-law slope with fixed , convolved with an exponential dust law with index and . This simple model reproduces the curvature of the continuum well over the considered wavelength range, better than a power-law of variable slope. The H line is modeled with a single broad-line Gaussian profile, while H has a higher SNR and is therefore more complicated, requiring a broad-line and a narrow-line component, even at the JWST/NIRSpec prism resolution (effectively at m for point sources, see de Graaff et al. 2023). Fainter lines adjacent to H (the [O III]Å-doublet, He IIÅ, and H) are masked. The [N II]Å-doublet is undetectable and including it in the fit changes the H by 1%, which is why we omit it. Finally, we include two nuisance parameters: One is a rescaling of the resolution of the spectrum by a factor with a uniform prior, reflecting uncertainty in the JWST/NIRSpec LSF. A value of is appropriate for uniformly exposed shutters and may be appropriate for point sources (de Graaff et al. 2023). The second nuisance parameter is with a logarithmic prior, reflecting a possible over- or underestimate of the noise in the fit. The posterior distributions are then sampled with PyMultiNest (Buchner et al. 2014).
@@
First, we fit the spectrum of image A (epoch 7), which has the highest SNR, with the broad- and narrow-line widths as free parameters, finding ,km/s, km/s, and . In the subsequent fitting of the spectra of images B, C, and A (epoch 8), we fix the line widths and LSF scaling to the values of image A to ensure consistency. Typical values of noise_scale are , indicating that the errors in the spectrum were only slightly underestimated. Finally, from the best-fit model posterior samples, we determine for H and H the continuum at the line center, the total line flux, and the EW, taking into account parameter correlations. For H, we sum the narrow and broad components as they are difficult to separate reliably. Basic convergence tests were done (live points, sampling efficiency), showing the measurements were robust. The resulting continuum fluxes, integrated line fluxes and EWs are listed in Tab. 3 in section 3.

### v3

@@ Acknowledgements.
We would like to thank Xihan Ji, Hannah Übler, and Roberto Maiolino, for cordial and useful discussions. The BGU lensing group acknowledges support by grant No. 2020750 from the United States-Israel Binational Science Foundation (BSF) and grant No. 2109066 from the United States National Science Foundation (NSF), and by the Israel Science Foundation Grant No. 864/23. P.D. warmly thanks the European Commission’s and University of Groningen’s CO-FUND Rosalind Franklin program. This work is based on observations obtained with the NASA/ESA/CSA JWST, namely programs GO-2756, -2561, -2883, -3538, -4111, and -3516, retrieved from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI). STScI is operated by the Association of Universities for Research in Astronomy, Inc. under NASA contract NAS 5-26555. The spectroscopy products presented herein, from JWST program GO-2561, were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. The data used in this work may be retrieved from the MAST archive at: http://dx.doi.org/10.17909/p7t7-te67. This work also makes use of the Center for Computational Astrophysics at the Flatiron Institute which is supported by the Simons Foundation. Support for JWST programs GO-2561, -4111, and -3516 was provided by NASA through grants from STScI. This research made use of Astropy,444http://www.astropy.org a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018) and Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2024), as well as the packages NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), and the MAAT Astronomy and Astrophysics tools for MATLAB (Ofek, 2014).References
@@
- Baggen et al. (2024) Baggen, J. F. W., van Dokkum, P., Brammer, G., et al. 2024, ApJ, 977, L13
- Baggen et al. (2025) Baggen, J. F. W., van Dokkum, P., Brammer, G., et al. 2025, ApJ, 980, L47
@@
- Bezanson et al. (2024) Bezanson, R., Labbe, I., Whitaker, K. E., et al. 2024, ApJ, 974, 92
@@
- de Graaff et al. (2024a) de Graaff, A., Brammer, G., Weibel, A., et al. 2024a, arXiv e-prints, arXiv:2409.05948
- de Graaff et al. (2024b) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024b, A&A, 684, A87
- de Graaff et al. (2025) de Graaff, A., Rix, H.-W., Naidu, R. P., et al. 2025, arXiv e-prints, arXiv:2503.16600
- D’Eugenio et al. (2025) D’Eugenio, F., Maiolino, R., Perna, M., et al. 2025, arXiv e-prints, arXiv:2503.11752
@@
- Euclid Collaboration et al. (2025) Euclid Collaboration, Bisigello, L., Rodighiero, G., et al. 2025, arXiv e-prints, arXiv:2503.15323
@@
- Golubchik et al. (2024) Golubchik, M., Steinhardt, C. L., Zitrin, A., et al. 2024, ApJ, 976, 108
@@
- Inayoshi & Maiolino (2025) Inayoshi, K. & Maiolino, R. 2025, ApJ, 980, L27
@@
- Killi et al. (2024) Killi, M., Watson, D., Brammer, G., et al. 2024, A&A, 691, A52
@@
- Li et al. (2025) Li, Z., Inayoshi, K., Chen, K., Ichikawa, K., & Ho, L. C. 2025, ApJ, 980, 36
- Lin et al. (2025) Lin, R., Zheng, Z.-Y., Jiang, C., et al. 2025, ApJ, 980, L34
@@
- Ma et al. (2025) Ma, Y., Greene, J. E., Setton, D. J., et al. 2025, ApJ, 981, 191
@@
- Maiolino et al. (2025) Maiolino, R., Risaliti, G., Signorini, M., et al. 2025, MNRAS, 538, 1921
@@
- Naidu et al. (2025) Naidu, R. P., Matthee, J., Katz, H., et al. 2025, arXiv e-prints, arXiv:2503.16596
@@
- Perger et al. (2025) Perger, K., Fogasy, J., Frey, S., & Gabányi, K. É. 2025, A&A, 693, L2
- Price et al. (2025) Price, S. H., Bezanson, R., Labbe, I., et al. 2025, ApJ, 982, 51
@@
- Rusakov et al. (2025) Rusakov, V., Watson, D., Nikopoulos, G. P., et al. 2025, arXiv e-prints, arXiv:2503.16595
- Schindler et al. (2024) Schindler, J.-T., Hennawi, J. F., Davies, F. B., et al. 2024, arXiv e-prints, arXiv:2411.11534
@@
- Setton et al. (2025) Setton, D. J., Greene, J. E., Spilker, J. S., et al. 2025, arXiv e-prints, arXiv:2503.02059
- Suess et al. (2024) Suess, K. A., Weaver, J. R., Price, S. H., et al. 2024, ApJ, 976, 101
- Tee et al. (2025) Tee, W. L., Fan, X., Wang, F., & Yang, J. 2025, ApJ, 983, L26
- Treu et al. (2022) Treu, T., Roberts-Borsani, G., Bradac, M., et al. 2022, ApJ, 935, 110
- Tripodi et al. (2024) Tripodi, R., Martis, N., Markov, V., et al. 2024, arXiv e-prints, arXiv:2412.04983
@@
- Yue et al. (2024) Yue, M., Eilers, A.-C., Ananna, T. T., et al. 2024, ApJ, 974, L26
@@ Appendix A Photometric measurements and systematic uncertainties
– Note: A summary of epochs is given in Tab. 2. Note that the uncertainties here do not contain systematics such as magnification uncertainties or the systematic scatter between epochs. The filters listed here are only those used in section 4, i.e. the bands that have coverage in at least one epoch in 2022 and 2023 each. The NIRCam epoch 5, listed in Tab. 2, only contains medium bands that are not observed in any other epoch and is therefore not used further in section 4.
@@ Appendix B Spectroscopic measurements
For spectroscopy, we use the latest reductions of the UNCOVER spectroscopic data (Price et al. 2025) available from the DJA (Heintz et al. 2024; de Graaff et al. 2024a). The spectra are reduced with MSAEXP v0.9.5 and reach m, fully covering H at . The spectra of images A and B have typical per pixel in the continuum, while the spectrum of images C, and A in epoch 8, have . In addition to the uncertainties provided by MSAEXP, we adopt a systematic error floor of 3 % per pixel in the continuum and 6 % per pixel at the location of strong lines. These errors are added in quadrature before fitting the spectra to account for spectro-photometric calibration uncertainties and other uncertainties related to the detailed shape of the intrinsic lines and the NIRSpec line-spread-function (LSF). The spectra are modeled over an observed wavelength-range m, i.e. red-ward of the red-shifted Balmer break, as we are here primarily interested in a robust measurement of the strongest Balmer lines and the continuum at the location of said lines.
@@
The model fitted to the spectra consists of the following components: The continuum is modeled with a power-law slope , convolved with an exponential dust law with index and . This simple model reproduces the curvature of the continuum well over the considered wavelength range, better than a power-law of variable slope. Note that fitting with the latter only changes the results of H and H by %. The H line is modeled with a single broad-line Gaussian profile since the spectral resolution at the observed wavelength () is not sufficient to resolve the narrow component, as was already shown in Furtak et al. (2024). H on the other hand has both a higher SNR and a higher spectral resolution. It is therefore more complicated, requiring a broad-line and a narrow-line component, even at the JWST/NIRSpec prism resolution (effectively at m for point sources, see de Graaff et al. 2024b). Fainter lines adjacent to H (the [O III]Å-doublet, He IIÅ, and H) are masked. The [N II]Å-doublet is undetectable and including it in the fit changes the H by 1%, which is why we omit it. The model is convolved with the wavelength-dependent NIRSpec prism LSF for a point-source (de Graaff et al. 2024b). Finally, we include two nuisance parameters: One is a rescaling of the resolution of the spectrum by a factor with a uniform prior, reflecting uncertainty in the JWST/NIRSpec LSF. A value of is appropriate for uniformly exposed shutters and may be appropriate for point sources (de Graaff et al. 2024b). The second nuisance parameter is with a logarithmic prior, reflecting a possible over- or underestimate of the noise in the fit. We list all free parameters and their priors in Tab. 5. The posterior distributions are then sampled with PyMultiNest (Buchner et al. 2014).
@@
First, we fit the spectrum of image A (epoch 7), which has the highest SNR, with the broad- and narrow-line widths as free parameters, finding km/s for both broad lines, km/s, and . In the subsequent fitting of the spectra of images B, C, and A (epoch 8), we fix the line widths and LSF scaling to the values of image A to ensure consistency. Typical values of noise_scale are , indicating that the errors in the spectrum were only slightly underestimated. Finally, from the best-fit model posterior samples, we determine for H and H the continuum at the line center, the total line flux, and the EW, taking into account parameter correlations. For H, we sum the narrow and broad components as they are difficult to separate reliably. Basic convergence tests were done (live points, sampling efficiency), showing the measurements were robust. The resulting continuum fluxes, integrated line fluxes and EWs are listed in Tab. 3 in section 3.
| Parameter | Prior type | Prior bounds |
| uniform | ||
| uniform | ||
| uniform | ||
| log-uniform | ||
| [km/s] | uniform | |
| [km/s] | uniform | |
| log-uniform | ||
| log-uniform | ||
| log-uniform | ||
| uniform | ||
| log-uniform |
– Note: The spectra are forward-modeled in rest-frame to reduce uncertainties.